# Statistical modeling to adjust for time trends in adaptive platform trials utilizing non-concurrent controls

### v1

@@ Abstract
Utilizing non-concurrent controls in the analysis of late-entering experimental arms in platform trials has recently received considerable attention, both on academic and regulatory levels. While incorporating this data can lead to increased power and lower required sample sizes, it might also introduce bias to the effect estimators if temporal drifts are present in the trial. Aiming to mitigate the potential calendar time bias, we propose various frequentist model-based approaches that leverage the non-concurrent control data, while adjusting for time trends. One of the currently available frequentist models incorporates time as a categorical fixed effect, separating the duration of the trial into periods, defined as time intervals bounded by any treatment arm entering or leaving the platform. In this work, we propose two extensions of this model. First, we consider an alternative definition of the time covariate by dividing the trial into fixed-length calendar time intervals. Second, we propose alternative methods to adjust for time trends. In particular, we investigate adjusting for autocorrelated random effects to account for dependency between closer time intervals and employing spline regression to model time with a smooth polynomial function. We evaluate the performance of the proposed approaches in a simulation study and illustrate their use by means of a case study.
@@ 1 Introduction
Recently, several modeling approaches were proposed to overcome the challenge of leveraging the NCC data, while still controlling the type I error rate and without introducing bias to the effect estimators. Focusing on a platform trial with two experimental arms, Lee and Wason [8] and Bofill Roig et al. [9] considered frequentist modeling approaches, where time is included to the model either as a linear function, or a categorical covariate. For the latter, the trial is split into two periods, separated by the time point where the second experimental arm was added to the platform. It was shown that both models improve the statistical power as compared to the separate analysis and control the type I error rate, if the functional form of the time trend is correctly specified, the time trends are equal across all arms and additive on the model scale [9]. Moreover, the model with categorical adjustment asymptotically maintains the type I error rate even under misspecification of the functional form of the time trend, provided that block randomization is used. To deal with situations where time trends could differ between arms, Bofill Roig et al. [9] also investigated a model that includes interaction between treatment and time. However, while this model maintains the type I error rate also in cases with different time trends, the power gain from using NCC data is lost. A Bayesian alternative was suggested by Saville et al. [10] in their so-called Time Machine approach, which uses a Bayesian hierarchical model and adjusts for time in terms of intervals (“buckets”) of equal length. The model smooths over the control response across the time buckets, such that closer buckets are modeled with more similar responses. They show that the Time Machine approximately controls the type I error rate and leads to superior performance in terms of statistical power as compared to the frequentist model with categorical adjustment for time in scenarios with linear time trend pattern. Marschner and Schou [12] proposed to analyze platform trials using meta-analysis techniques, which allow to conduct both treatment-control and treatment-treatment comparisons for non-concurrent arms. In this network meta-analysis approach, the platform trial is viewed as a network of direct (concurrent) comparisons, made within periods between trial adaptations; and indirect (non-concurrent) comparisons, made across multiple stages in the trial using a common reference arm. The estimates of the non-concurrent treatment-control comparisons are then obtained by linearly combining the direct contrast estimates.
@@
In this work, we extend the frequentist modeling strategies from Bofill Roig et al. [9]. In particular, we propose an alternative definition of the time variable and consider adjusting for equally spaced calendar time intervals, similarly to the Bayesian Time Machine, also in the frequentist setting. Moreover, we investigate more advanced modeling approaches, such as mixed models or polynomial splines. The assumption of equal time trends, required for all methods proposed so far, may not always hold in practice, as temporal changes may affect treatment arms differently, e.g. in settings with different disease variants. Therefore, we investigate whether this assumption can be relaxed by considering a mixed model, which includes the interaction between treatment and time as a random effect. The performance of the proposed methods is evaluated in a simulation study, where we also assess the conditions under which the approaches lead to valid statistical inference. The use of the considered methods is demonstrated on a real-world dataset from a clinical trial for a rare disease - progressive supranuclear palsy.
@@
The remainder of this paper is structured as follows: Section 2 introduces the notation and considered trial design and describes the proposed methods for platform trials with equal time trends across all arms. In Section 3 we discuss a possible extension to trials where the temporal changes affect the arms differently. Section 4 presents an extensive simulation study, evaluating the operating characteristics of the considered methods under a wide range of settings. The use of the modeling strategies is exemplified in a case study using data on a rare neurodegenerative disease, presented in Section 5. We conclude the paper with a discussion and conclusions in Section 6.
@@ 2 Methods
To fit the proposed models, we use all data from the trial until treatment $M$ leaves the platform. In other words, we use all data in the set $\mathcal{D}_{M}=\{(y_{j},k_{j},t_{j}),j=1,...,N\mid t_{j}\leq t^{exit}_{M}\}$, where $j$ represents the patient index, $y_{j}$ denotes the response, $k_{j}$ and $t_{j}$ are the arm and recruitment time corresponding to patient $j$, respectively, and $t^{exit}_{M}$ marks the time point when the arm $M$ left the trial. Note that also data from arms for which the recruitment is still ongoing at time $t^{exit}_{M}$ is used to fit the models.
For technical definitions of the time variables, periods, and calendar time intervals, we refer to the supplementary material (see Sect. A).
@@ 2.1 Fixed effect models
Period adjustment. As previously investigated by Lee and Wason [8] and Bofill Roig et al. [9] in simple platform trials with two experimental arms only, in the first model, we adjust for the time effect by including the factor “period” to the model, i.e., using a step-function:
@@
Here $\eta_{0,calendar}$ represents the control response in the first calendar time unit; $\theta_{k}$ denotes the treatment effect of treatment $k$ compared to control for $k\in\mathcal{K}_{M}$; $\tau_{c}$ is the stepwise effect between calendar time units 1 and $c$ ($c=2,\ldots,C_{M}$), where $C_{M}$ indicates the calendar time unit, in which arm $M$ left the trial. The length of the calendar time units $c_{length}$ is given in terms of the number of enrolled patients and can be specified arbitrarily.
Note that both models (1) and (2) implicitly assume a time trend that is constant in all periods or calendar time intervals, respectively, and that the time effect is equal across all arms and additive on the model scale. In both models, the residuals $\varepsilon_{j}$ are assumed to follow a normal distribution $\varepsilon_{j}\sim\mathcal{N}(0,\sigma^{2})$. However, the models take slightly different datasets into account - model (1) all trial data until the period $S_{M}$, model (2) all data until calendar unit $C_{M}$, which typically do not end at the same time. Moreover, the length of the time intervals in (2) poses an additional design parameter that allows to adjust for shorter time intervals than those given by the periods, which are also independent of the alternations to the trial design.
@@ 2.2 Spline regression
Specifically, we consider the B-spline function to model the time trend. This function is composed of multiple polynomial functions of a given degree $q$, which are joined together at points called knots, such that the entire spline is continuously differentiable up to the $(q-1)$th derivative [13]. For the positions of these knots, we evaluate two strategies, somewhat analogous to the period and calendar time adjustments. Firstly, we place the inner knots to the beginning of each period $s=2,\ldots,S$, such that one polynomial of degree $q$ is always fitted to each period. Moreover, we consider placing the inner knots equidistantly, according to the length of the calendar time unit, and thus fitting a polynomial of degree $q$ to each calendar time interval. Regarding the degree of the fitted polynomial, we explore linear, quadratic and cubic splines, i.e. $q\in(1,2,3)$. For details on the definition of the B-spline, see Sect. B.1 in the supplementary material.
Modeling the time using spline functions gives the model additional flexibility compared to the regression model that adjusts for time using a linear function considered in Lee and Wason [8] and Bofill Roig et al. [9], as now also more complex time trend patterns can be modeled more accurately.
@@ 4.1.1 Data generation
[…]
@@
  • Seasonal time trend: $f(j)=\lambda_{k}\cdot\mathrm{sin}\big{(}\psi\cdot 2\pi\cdot\frac{j-1}{N-1}\big{)}$, where $N$ indicates the total sample size in the trial. The seasonal trend may consist of multiple cycles, where the response increases at first and decreases afterwards, while the respective peaks of the cycles correspond to $\lambda_{k}$ or $-\lambda_{k}$. The number of cycles over the whole platform trial is determined by $\psi$.
@@ 4.1.2 Considered trial settings
Setting 2: In order to understand the behavior of the fixed effect model using the alternative definition of time, as well as the mixed effect models, it is sufficient to investigate a simpler platform trial with $K=4$ experimental arms and equal time trend in all arms ($\lambda_{k}=\lambda_{0}$ $\forall k$). We compare arm $M=3$ to the control, while varying the pattern (considering all patterns listed above) and strength of the time trend ($\lambda_{k}$).
@@
Setting 3: Keeping the platform trial with $K=4$ experimental arms from the previous setting, we focus on examining trials with different time trends across arms. Again, we compare arm $M=3$ to the control, while varying the pattern (considering all patterns listed above) and strength of the time trend ($\lambda_{k}$). In particular, the treatment arm 3 and the control group have no time trend ($\lambda_{0}=\lambda_{3}=0$), while the rest of the arms have time trends of different strengths. We vary the strength of the time trend in arms 1 ($\lambda_{0}=\lambda_{2}=\lambda_{3}=\lambda_{4}\neq\lambda_{1}$), 1 and 2 ($\lambda_{0}=\lambda_{3}=\lambda_{4}\neq\lambda_{1}=\lambda_{2}$), or 1, 2 and 4 ($\lambda_{0}=\lambda_{3}\neq\lambda_{1}=\lambda_{2}=\lambda_{4}$ and $\lambda_{0}=\lambda_{3}\neq\lambda_{1}$; $\lambda_{2}=2\lambda_{1}$; $\lambda_{4}=3\lambda_{1}$). In this setting, we evaluate the performance of the mixed effect models with interaction ((6) and (7)) and compare it to the fixed effect model adjusting for periods (1).
@@
4) $\lambda_{1}\in[-0.5,0.5];\leavevmode\nobreak\ \lambda_{2}=2\lambda_{1};\leavevmode\nobreak\ \lambda_{4}=3\lambda_{1}$;
@@
Table 1: Simulation settings and parameters considered in the simulation study.
@@ 4.2 Results
Setting 1A. Figure 4 shows the impact of the strength of the time trend on the operating characteristics when evaluating treatment arm 5. Analogous plots for all treatment arms can be found in Figure S1 in the supplementary material. The regression model, as well as the separate analysis asymptotically control the type I error rate, regardless of the strength of the time trend. The pooled analysis leads to inflation of the type I error in the presence of positive time trends and deflation in case of negative time trends. Although only results for linear time trends are shown, the regression model maintains the type I error rate under arbitrary time trend pattern. Additionally, the model leads to gains in power as compared to the separate analysis. This result is in line with previous works [8, 9].
@@
We vary the overlap between arms by varying the parameter $d$ from 0 to $5002n=$ 500. The effect of the amount of overlapping sample size when evaluating treatment arm 5 is shown in Figure 5. Figure S2 in the supplementary material shows analogous plots for all treatment arms. The overlaps do not affect the type I error rate control of the regression model and separate approach, which is guaranteed (asymptotically) in all the cases. The inflation of the pooled analysis gets stronger with increasing $d$. This is because larger values of $d$ result in longer platform trials and larger size of the NCC data. The power of the regression model, however, depends on the overlap between treatment arms. In the extreme cases with $d=0$ and $d=2n$, the regression model leads to identical power as the separate analysis. If $d=0$, there is no NCC data, as all the arms join the trial at the beginning. Thus, the control group used for the treatment-control comparison is the same for both, the separate analysis and the regression model. In case of no overlap between the arms ($d=2n=500$), there is insufficient amount of data to estimate the period effect. Hence, simultaneous presence of the experimental arms in the trial is crucial for a reliable estimation of the period effect and resulting power gains when using the regression model. For the considered simulations, the maximal power is reached for $d=175$, which in this case leads to the optimal trade-off between the amount of overlapping sample size between treatment arms and the size of the non-concurrent data.
@@
Figure S3 in the supplementary material illustrates how the type I error rate and power for individual treatment-control comparisons depends on the order of entry in the platform trial. We observe that the power of the regression model increases for arms that were added to the trial later. This is due to larger sample size of the NCC data. On the other hand, this also leads to a higher inflation of the type I error with the pooled analysis.
@@
Setting 1B. To evaluate the performance of the spline regression, we focused on a platform trial where arm $k$ is added after every $d_{k}=250\cdot(k-1)$ patients have been enrolled in the trial. Here, we also varied the degree of the B-splines in the simulations, considering linear, quadratic, and cubic splines. However, since the differences in the resulting operating characteristics were only marginal, here we only present results for the cubic spline regression. The knots are placed either at the beginning of each period or each calendar time interval. Results for linear and quadratic splines can be found in supplementary Figure S4.
@@
The type I error rate and power of the evaluated models under the considered scenarios when comparing treatment arm 5 to the control are presented in Figure 6. Analogous plots for other experimental arms are given in Figure S5 in the supplement. If the time trend pattern is given by a continuous function, the spline regression maintains the type I error rate. However, in case of the stepwise time trend, we observe an inflation in the type I error rate, particularly pronounced when placing the knots according to the periods. This is because the spline regression estimates the time effect by a smooth function, composed of multiple polynomial functions joined together in the knots. This is not an optimal approach if there are sudden jumps in the time trend. Note that the separate approach has a type I error rate strictly below 2.5% for strong time trends under the seasonal trend pattern and in particular under the stepwise trend pattern, as this one has stronger values than the other shapes. The reason is that block randomization was used to assign the patients to the active arms. In case of very strong time trends, ignoring the blocks in the analysis results in a conservative test, as the variances of the treatment effects are overestimated. This leads to deflation in the type I error rate, which gets stronger the more extreme the time trend is [16].
The power of the spline regression in cases with continuous time trend function is slightly improved as compared to the fixed effects model, however, the differences are only marginal in the considered scenarios.
@@
Figure 6: Setting 1B: Type I error rate and power for the cubic spline regression model with knots according to periods or calendar time units compared to the regression model with period adjustment with respect to the strength of the time trend $\lambda$ using different time trend patterns. Here, $n=250$ in each experimental arm and $N_{p}=1660$ in case of inverted-U trend (corresponding to the middle of the trial) are used. In case of calendar time adjustment, unit size of 100 patients is considered. Treatment arm 5 is being evaluated.

### v2

@@ Abstract
Utilizing non-concurrent control data (NCC) in the analysis of late-entering arms in platform trials has recently received considerable attention. While incorporating NCC can lead to increased power and lower sample sizes, it might introduce bias to the effect estimators if temporal drifts are present. Aiming to mitigate this potential bias, we propose various frequentist model-based approaches that leverage the NCC, while adjusting for time. One of the currently available models incorporates time as a categorical fixed effect, separating the trial duration into periods, defined as time intervals bounded by any arm entering or leaving the platform. In this work, we propose two extensions of this model. First, we consider an alternative definition of time by dividing the trial into fixed-length calendar time intervals. Second, we propose alternative model-based time adjustments. Specifically, we investigate adjusting for random effects and employing splines to model time with a polynomial function. We evaluate the performance of the proposed approaches in a simulation study and illustrate their use through a case study. We show that adjusting for time via a spline function controls the type I error in trials with a sufficiently smooth time trend pattern and may lead to power gains compared to the standard fixed effect model. However, the fixed effect model with period adjustment is the most robust model for arbitrary time trends, provided that the trend is equal across all arms. Especially, in trials with sudden changes in the time trend, the period-adjustment model is preferred if NCC are included.
@@ 1 Introduction
Recently, several modeling approaches were proposed to overcome the challenge of leveraging the NCC data, while still controlling the type I error rate and without introducing bias to the effect estimators. Focusing on a platform trial with two experimental arms, [8] and [9] considered frequentist modeling approaches, where time is included to the model either as a linear function, or a categorical covariate. For the latter, the trial is split into two periods, separated by the time point where the second experimental arm was added to the platform. It was shown that both models improve the statistical power as compared to analysis using CC data only and control the type I error rate, if the functional form of the time trend is correctly specified, the time trends are equal across all arms and additive on the model scale [9]. Moreover, the model with categorical adjustment asymptotically maintains the type I error rate even under misspecification of the functional form of the time trend, provided that block randomization is used. To deal with situations where time trends could differ between arms, [9] also investigated a model that includes interaction between treatment and time. However, while this model maintains the type I error rate also in cases with different time trends, the power gain from using NCC data is lost. A Bayesian alternative was suggested by [10] in their so-called Time Machine approach, which uses a Bayesian hierarchical model and adjusts for time in terms of intervals (“buckets”) of equal length. The model smooths over the control response across the time buckets, such that closer buckets are modeled with more similar responses. They show that the Time Machine approximately controls the type I error rate and leads to superior performance in terms of statistical power as compared to the frequentist model with categorical adjustment for time in scenarios with linear time trend pattern. [12] proposed to analyze platform trials using meta-analysis techniques, which allow to conduct both treatment-control and treatment-treatment comparisons for arms that are not simultaneously active in the trial. In this network meta-analysis approach, the platform trial is viewed as a network of direct (concurrent) comparisons, made within periods between trial adaptations; and indirect (non-concurrent) comparisons, made across multiple stages in the trial using a common reference arm. The estimates of the non-concurrent treatment-control comparisons are then obtained by linearly combining the direct contrast estimates.
@@
In this work, we extend the frequentist modeling strategies from [9]. In particular, we propose an alternative definition of the time variable and consider adjusting for equally spaced calendar time intervals, similarly to the Bayesian Time Machine, also in the frequentist setting. Moreover, we investigate more advanced modeling approaches, such as mixed models or polynomial splines. The assumption of equal time trends, required for all methods proposed so far, may not always hold in practice, as temporal changes may affect treatment arms differently, e.g. in settings with different disease variants. Therefore, we investigate whether this assumption can be relaxed by considering a mixed model, which includes the interaction between treatment and time as a random effect. The performance of the proposed methods is evaluated in a simulation study, where we also assess the conditions under which the approaches lead to valid statistical inference. The use of the considered methods is demonstrated on a real-world dataset from a clinical trial for a rare disease (progressive supranuclear palsy), and in a hypothetical study based on a real platform trial for chronic lymphocytic leukemia (presented in supplementary material, Sect. D).
@@
The remainder of this paper is structured as follows: Section 2 introduces the notation and considered trial design and describes the proposed methods for platform trials with equal time trends across all arms. In Section 3, we discuss a possible extension to trials where the temporal changes affect the arms differently. Section 4 presents an extensive simulation study, evaluating the operating characteristics of the considered methods under a wide range of settings. The use of the modeling strategies is exemplified in two case studies using data on a rare neurodegenerative disease and a type of blood cancer, presented in Section 5. We conclude the paper with a discussion and conclusions in Section 6.
@@ 2 Methods
To fit the proposed models, we use all data from the trial until treatment $M$ leaves the platform. In other words, we use all data in the set $\mathcal{D}_{M}=\{(y_{j},k_{j},t_{j}),j=1,...,N\mid t_{j}\leq t^{exit}_{M}\}$, where $j$ represents the patient index, $y_{j}$ denotes the response, $k_{j}$ and $t_{j}$ are the arm and recruitment time corresponding to patient $j$, respectively, and $t^{exit}_{M}$ marks the time point when the arm $M$ left the trial. Note that also data from arms for which the recruitment is still ongoing at time $t^{exit}_{M}$ is used to fit the models. For technical definitions of the time variables, periods, and calendar time intervals, we refer to the supplementary material (see Sect. A).
@@ 2.1 Fixed effect models
Period adjustment. As previously investigated by [8] and [9] in simple platform trials with two experimental arms only, in the first model, we adjust for the time effect by including the factor “period” to the model, i.e., using a step-function:
@@
Here $\eta_{0,calendar}$ represents the control response in the first calendar time unit; $\theta_{k}$ denotes the treatment effect of treatment $k$ compared to control for $k\in\mathcal{K}_{M}$; $\tau_{c}$ is the stepwise effect between calendar time units 1 and $c$ ($c=2,\ldots,C_{M}$), where $C_{M}$ indicates the calendar time unit, in which arm $M$ left the trial. The length of the calendar time units $c_{length}$ is given in terms of the number of enrolled patients and can be specified arbitrarily. Note that since only data until $t_{M}^{exit}$ are used to fit the model, the last calendar time unit $C_{M}$ will typically be shorter than $c_{length}$, as entering and leaving times of the experimental treatment arms are independent of the calendar time unit length $c_{length}$.
Both models (1) and (2) implicitly assume a time trend that is constant in all periods or calendar time intervals, respectively, and that the time effect is equal across all arms and additive on the model scale. In both models, the residuals $\varepsilon_{j}$ are assumed to follow a normal distribution $\varepsilon_{j}\sim\mathcal{N}(0,\sigma^{2})$. Moreover, the length of the time intervals in (2) poses an additional design parameter that allows to adjust for longer or shorter time intervals than those given by the periods, which are also independent of the alternations to the trial design.
@@ 2.2 Spline regression
Specifically, we consider the B-spline function to model the time trend. This function is composed of multiple polynomial functions of a given degree $q$, which are joined together at points called knots, such that the entire spline is continuously differentiable up to the (q1)(q-1)th derivative [13]. For the positions of these knots, we evaluate two strategies, somewhat analogous to the period and calendar time adjustments. Firstly, we place the inner knots to the beginning of each period $s=2,\ldots,S$, such that one polynomial of degree $q$ is always fitted to each period. Moreover, we consider placing the inner knots equidistantly, according to the length of the calendar time unit, and thus fitting a polynomial of degree $q$ to each calendar time interval. Regarding the degree of the fitted polynomial, we explore linear, quadratic and cubic splines, i.e. $q\in(1,2,3)$. For details on the definition of the B-spline, see Sect. B.1 in the supplementary material.
Modeling the time using spline functions gives the model additional flexibility compared to the regression model that adjusts for time using a linear function considered in [8] and [9], as it allows for more accurate modeling of complex time trend patterns.
@@ 4.1.1 Data generation
To generate the trial data we assume an allocation ratio of $1:1:\ldots:1$ in each period. Patients are assigned to arms following block randomization with block sizes of $2\cdot(\text{number of active arms}+1)$ at each time $t$. Patients are indexed by entry order, assuming that at each time unit exactly one patient is recruited (i.e., $t_{j}=j$) and the time of recruitment and of the observation of the response are equal. The continuous outcome $y_{j}$ for patient $j$ is then drawn from a normal distribution according to $y_{j}\sim\mathcal{N}(\mu,\sigma^{2})$ with
@@
  • Seasonal time trend: $f(j)=\lambda_{k}\cdot\mathrm{sin}\big{(}\psi\cdot 2\pi\cdot\frac{j-1}{N-1}\big{)}$, where $N$ indicates the total sample size in the trial. The seasonal trend may consist of multiple cycles, where the response increases at first and decreases afterwards, while the respective peaks of the cycles correspond to $\lambda_{k}$ or $\lambda_{k}$. The number of cycles over the whole platform trial is determined by $\psi$.
@@ 4.1.2 Considered trial settings
Setting 2: In order to understand the behavior of the fixed effect model using the alternative definition of time, as well as the mixed effect models, we investigate a simpler platform trial with $K=4$ experimental arms and equal time trend in all arms ($\lambda_{k}=\lambda_{0}$ $\forall k$). We compare arm $M=3$ to the control, while varying the pattern (considering all patterns listed above) and strength of the time trend ($\lambda_{k}$).
@@
Setting 3: Keeping the platform trial with $K=4$ experimental arms from the previous setting, we focus on examining trials with different time trends across arms. Again, we compare arm $M=3$ to the control, while varying the pattern (considering all patterns listed above) and strength of the time trend ($\lambda_{k}$). In particular, the treatment arm 3 and the control group have no time trend ($\lambda_{0}=\lambda_{3}=0$), while the rest of the arms have time trends of different strengths. We vary the strength of the time trend in arms 1 ($\lambda_{0}=\lambda_{2}=\lambda_{3}=\lambda_{4}\neq\lambda_{1}$), 1 and 2 ($\lambda_{0}=\lambda_{3}=\lambda_{4}\neq\lambda_{1}=\lambda_{2}$), or 1, 2 and 4 ($\lambda_{0}=\lambda_{3}\neq\lambda_{1}=\lambda_{2}=\lambda_{4}$ and $\lambda_{0}=\lambda_{3}\neq\lambda_{1}$; $\lambda_{2}=2\lambda_{1}$; $\lambda_{4}=3\lambda_{1}$). In this setting, we evaluate the performance of the mixed effect models with interaction ((6) and (7)) and compare it to the fixed effect model adjusting for periods (1). Note that we only consider cases where for the evaluated treatment arm, the time trend is equal to the one in the control group. If this assumption does not hold, the considered testing approaches would test the null hypothesis that the average treatment effect over the time when the treatment is active in the platform is lower or equal to 0. This would imply that the treatment effect depends on the time periods, in which the treatment is active in the platform.
@@
Main findings in the
considered settings
Conditions
for T1E control
@@
The fixed effect model with period adjustment
leads to power gains compared to the separate
analysis, provided that there is some overlap
between the treatment arms.
Equal and additive
time trends, block
randomization.
@@
Spline regression controls the type I error rate
if the time trend pattern is sufficiently smooth
and equal across all arms. In cases with little
to no overlap between the treatment arms,
splines achieve larger power than the regression
model with period adjustment.
Equal and additive
time trends, sufficiently
smooth trend pattern,
block randomization.
@@
Type I error rate inflation depends
on the unit size, and pattern and strength
of the time trend. Exact type I error rate
control is only achieved if calendar time
intervals match the entry and leaving times
of the experimental arms and the time trends
are equal across arms. Larger calendar
time units lead to larger power gains.
Equal and additive
time trends, appropriate
cal. time unit size,
block randomization.
@@
Mixed models with period as random factor
do not control the type I error rate in the
presence of time trends. Type I error rate
inflation is more pronounced in the case
of moderate time trends.
Type I error rate is not
controlled under positive
time trends and conservative
under negative time trends.
@@
4) $\lambda_{1}\in[-0.5,0.5]$
$\lambda_{2}=2\lambda_{1}$
$\lambda_{4}=3\lambda_{1}$
@@
Mixed models with fixed effect period and
the interaction between period and treatment
as random effect reduce the type I error rate
inflation in settings with unequal time trends
across arms compared to the fixed effect model
with period adjustment, but do not eliminate
it completely.
Equal and additive
time trends, appropriate
cal. time unit size,
block randomization.
Under unequal time trends
smaller T1E inflation than
fixed effect model (1).
Table 1: Simulation settings and parameters considered in the simulation study, and main findings for each considered scenario.
4.1.3 Implementation
The simulation study was performed using the NCC R package [16], using the version labeled as Release 1.4 on GitHub. The package contains functions implementing all considered analysis approaches, as well as functions for data simulation and performing simulation studies [17].
@@ 4.2 Results
Setting 1A. Figure 4 shows the impact of the strength of the time trend on the operating characteristics when evaluating treatment arm 5. Analogous plots for all treatment arms can be found in Figure S1 in the supplementary material. The regression model and the separate analysis asymptotically control the type I error rate, regardless of the strength of the time trend. The pooled analysis leads to inflation of the type I error in the presence of positive time trends and deflation in case of negative time trends. Additionally, the model leads to gains in power as compared to the separate analysis. This result is in line with previous works [8, 9], where the regression model was investigated for trials with two experimental arms only.
@@
We vary the overlap between arms by varying the parameter $d$ from 0 to $2n=500$. The effect of the amount of overlapping sample size when evaluating treatment arm 5 is shown in Figure 5. Figure S2 in the supplementary material shows analogous plots for all treatment arms. The overlaps do not affect the type I error rate control of the regression model and separate approach, which is guaranteed (asymptotically) in all the cases. The inflation of the pooled analysis gets stronger with increasing $d$. This is because larger values of $d$ result in longer platform trials and larger size of the NCC data. The power of the regression model, however, depends on the overlap between treatment arms. In the extreme cases with $d=0$ and $d=2n$, the regression model leads to identical power as the separate analysis. If $d=0$, there is no NCC data, as all the arms join the trial at the beginning. Thus, the control group used for the treatment-control comparison is the same for both, the separate analysis and the regression model. In case of no overlap between the arms ($d=2n=500$), there is insufficient amount of data to estimate the period effect. Hence, simultaneous presence of the experimental arms in the trial is crucial for a reliable estimation of the period effect and resulting power gains when using the regression model. For the considered simulations, the maximal power is reached for $d=175$, which in this case leads to the optimal trade-off between the amount of overlapping sample size between treatment arms and the size of the non-concurrent data.
@@
Figure S3 in the supplementary material illustrates how the type I error rate and power for individual treatment-control comparisons depends on the order of entry in the platform trial. We observe that the power of the regression model increases for arms that were added to the trial later. This is due to larger sample size of the NCC data. On the other hand, this also leads to a higher inflation of the type I error with the pooled analysis, while the regression model controls the type I error for all treatment-control comparisons.
To sum up, the fixed effect regression model with period adjustment offers superior performance in terms of the statistical power in trials where the treatment arms enter in a staggered way but overlap for some time. Arms added to the trial later have the advantage of more NCC data available for the treatment-control comparison, hence there is higher probability to identify efficacious experimental arms as such.
@@
Setting 1B. The performance of the spline regression was evaluated under varying overlap between the treatment arms, as well as time trends of different patterns and strengths. We also varied the degree of the B-splines in the simulations, considering linear, quadratic, and cubic splines. However, since the differences in the resulting operating characteristics were only marginal, here we only present results for the cubic spline regression. Results for linear and quadratic splines can be found in supplementary Figures S4 and S5. The knots are placed either at the beginning of each period or each calendar time interval.
@@
The type I error rate and power of the evaluated models under the considered scenarios when comparing treatment arm 5 to the control are presented in Figures 6 and 7. Analogous plots for other experimental arms are given in Figures S6 and S7 in the supplement. If the time trend pattern is given by a sufficiently smooth function, the spline regression maintains the type I error rate. However, in case of the stepwise time trend, we observe an inflation in the type I error rate. This is because the spline regression estimates the time effect by a smooth function, composed of multiple polynomial functions joined together in the knots. This is not a valid approach if there are sudden jumps in the time trend. The regression model maintains the type I error rate under an arbitrary time trend pattern and strength. Note that the separate approach has a type I error rate strictly below 2.5% for strong time trends under the seasonal trend pattern and in particular under the stepwise trend pattern, as this one has stronger values than the other shapes. The reason is that block randomization was used to assign the patients to the active arms. In case of very strong time trends, ignoring the blocks in the analysis results in a conservative test, as the variances of the treatment effects are overestimated. This leads to deflation in the type I error rate, which gets stronger the more extreme the time trend is [18].
In cases with less overlap between the treatment arms, the power of the spline regression is improved compared to the fixed effect model with period adjustment. In particular, in the case of no overlap ($d=500$), where the fixed effect model performs identically to the separate analysis, the spline regression achieves power gains of approximately 3 percentage points, while controlling the type I error rate for an arbitrary strength of the time trend, provided that its functional form is sufficiently smooth (see Figure 7).
@@
Figure 6: Setting 1B: Type I error rate and power for the cubic spline regression model with knots according to periods or calendar time units compared to the regression model with period adjustment with respect to the timing of adding the treatment arms using different time trend patterns. Here, $n=250$ in each experimental arm, $N_{p}$ placed in the middle of the trial in case of an inverted-U trend, and $\lambda=0.5$ are used. In case of calendar time adjustment, unit size of 450 patients is considered. Treatment arm 5 is being evaluated. In settings with inverted-U, linear, and seasonal time trends, the lines for splines with knots placed according to periods (pink) and calendar time units (blue) are nearly overlapping.
Figure 7: Setting 1B: Type I error rate and power for the cubic spline regression model with knots according to periods or calendar time units compared to the regression model with period adjustment with respect to the strength of the time trend $\lambda$ using different time trend patterns. Here, $n=250$ in each experimental arm, $d=500$ (no overlap between the arms), and $N_{p}=2500$ in case of inverted-U trend (corresponding to the middle of the trial) are used. In case of calendar time adjustment, unit size of 450 patients is considered. Treatment arm 5 is being evaluated. In the power plots, the lines for splines with knots placed according to periods (pink) and calendar time units (blue) are nearly overlapping, and the lines for fixed effect model with period adjustment (red) and separate analysis (black) are indistinguishable.